what is the donor concentration in n-type germanium of 1ohm-cm resistivity at 300K given electron mobility is equals to 3900cm²/V.s
1
Expert's answer
2019-05-06T10:34:15-0400
"ρ= \frac {1} {nqμ_n} (1)"

Using (1) we got:


"n=\frac {1} {ρqμ_n} (2),"

where ρ=1ohm×cm, q=1.6×10-19 C, μn=3900cm²/V.s

In this case, n=1.6×1014 cm-3


Answer:

1.6×1014 cm-3